Please help me with this multiple choice question...The parentheses stand for absolute value bars...Thanks!!
What is the solution set of the absolute value equation (x^2 - 2x)= 3x-6 ?
a) (2,not = 3) b) 2 c) not = 3 d)(2,3)

I'm not sure whether to set it equal to zero or what side of the equation to make negative...thank you!

Answer by nerdybill(6963)

(x^2 - 2x)= 3x-6
x^2 - 2x = 3x-6
x^2 - 5x = -6
x^2 - 5x + 6= 0
(x-2)(x-3) = 0
x = {2,3} (answer d)